OpenStudy (anonymous):

Is (7, 4) a solution of 3x - 4y = 5 yes or no

OpenStudy (unklerhaukus):

sub in x=7, and y=4, what does the equation become?

OpenStudy (anonymous):

True replace the equation \[3(7)-4(4)=5\] \[5=5\] So it is true

OpenStudy (muzzack):

well replace \[\Large\frak\ 3(7) - 4(4) = 5\] \[\Large\frak\ 14- 16 = 5\]

OpenStudy (muzzack):

not true

OpenStudy (anonymous):

\[21-16=5\] @Muzzack

OpenStudy (muzzack):

im sorry \[\Large\frak\ 21-16\]

OpenStudy (anonymous):

I'm going with yes lol Thanks guys.

OpenStudy (anonymous):

Anytime @Kemist and @Muzzack was on the right track.